For a segment of a radio​ show, a disc jockey can play 6 records. If there are 9 records to select​ from, in how many ways can the program for this segment be​ arranged?
Found 2 solutions by addingup, greenestamps:
Answer by addingup(3677)

When order doesn't matter:
9C6 = n!/(n-r)! = 9!/(9-6)! = 362880/6 = 60480

Answer by greenestamps(13200)



The response from the other tutor is confusing; he seems to be saying that order does not matter, but then his answer is if order DOES matter.

Since the question asks how many ways the program be arranged, to me that implies that order does matter. Then the number of ways is

P(9,6) = 9*8*7*6*5*4 = 60480.

If the question was intended only to ask which 6 of the 9 songs he could choose, with order not being important, then the number would be C(9,6) = C(9,3) = (9*8*7)/(3*2*1) = 84.